OpenStudy (anonymous):

Medal and Fan! Mrs. Parenti harvested tomatoes. She gave 1/4 of them to the Franks, 9 to the Fitzpatricks, and 5 to the Chins. She kept 4 for herself. How many tomatoes did she harvest?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

im patient. don't worry

OpenStudy (anonymous):

im sorry about this but i need ot go im trying ot find some one to hlep you

OpenStudy (anonymous):

ok don't sweat it :)

OpenStudy (anonymous):

18

OpenStudy (anonymous):

there is no 18 in the answer choises. only the following - 16 20 24 30

OpenStudy (anonymous):

well you can make an equation from that 1/4x+9+5+4=x move 1/4x to the other side 3/4x=9+5+4 combine like terms 3/4x=18 multiply by 4/3 x=18*4/3 x=72/3 x=24

OpenStudy (anonymous):

you could of have use the method of guess and check

OpenStudy (anonymous):

tnx first one to answer was the 1st one to get medal and an extra fan

OpenStudy (anonymous):

@Brolex1220 can you help me with my health question?